### v1

# PromptWise: Online Learning for Cost-Aware Prompt Assignment in Generative Models

Abstract
The rapid advancement of generative AI models has provided users with numerous options to address their prompts. When selecting a generative AI model for a given prompt, users should consider not only the performance of the chosen model but also its associated service cost. The principle guiding such consideration is to select the least expensive model among the available satisfactory options. However, existing model-selection approaches typically prioritize performance, overlooking pricing differences between models. In this paper, we introduce PromptWise, an online learning framework designed to assign a sequence of prompts to a group of large language models (LLMs) in a cost-effective manner. PromptWise strategically queries cheaper models first, progressing to more expensive options only if the lower-cost models fail to adequately address a given prompt. Through numerical experiments, we demonstrate PromptWise’s effectiveness across various tasks, including puzzles of varying complexity and code generation/translation tasks. The results highlight that PromptWise consistently outperforms cost-unaware baseline methods, emphasizing that directly assigning prompts to the most expensive models can lead to higher costs and potentially lower average performance.
1 Introduction
Recent advancements in large language models (LLMs) and text-guided generative AI have significantly broadened public access to various generative AI services (Rombach et al., 2022; Touvron et al., 2023; OpenAI et al., 2024; DeepSeek-AI et al., 2025). Users now have the opportunity to utilize and select among many well-functioning models developed by leading technology companies and research institutions. This extensive selection raises a fundamental question: how should users effectively select generative AI models to address their prompts?
A common approach is to evaluate each model based on expected performance scores and select the top-performing model to handle all prompts. However, as noted by recent studies, this strategy neglects the variability in model performance across different prompt categories. For example, as discussed in Qin et al. (2024); Frick et al. (2025); Hu et al. (2025), in comparing LLMs and text-to-image models, one candidate model could excel in responding to specific categories of prompts but perform suboptimally with the other prompt types. Such discrepancies could arise due to differences in training data and architectural choices. To perform a prompt-aware model selection, the recent literature has developed offline-trained selector models Qin et al. (2024); Frick et al. (2025) and online bandit algorithms Hu et al. (2025) aimed at adaptively assigning prompts to the prompt-specific optimal models in real-time.
Figure 1: Cost-aware prompt assignment example for the Sudoku-solving task: Our proposed PromptWise method can adaptively learn to assign easier Sudokus to the less expensive GPT-4o-mini model, alternating to the more expensive o3-mini model only for hard instances.
Although these existing offline and online learning methodologies considerably enhance overall prompt-handling performance, they commonly overlook the critical factor of the cost associated with each generative AI service. In practical scenarios, users are not constrained to assign prompts exclusively to a single model; they may utilize multiple models as long as the cumulative expense is economically justified. Consider, for example, a scenario where users handle easy and hard math questions using two LLMs (Figure 1): o3-mini (costing \$$4.4$ per million output tokens (PMT)) and GPT-4o-mini (costing \$$0.6$ PMT) [^1] https://platform.openai.com/docs/pricing. Given that o3-mini considerably outperforms GPT-4o-mini on complex tasks, a human-like assignment strategy involves directing challenging prompts to o3-mini while utilizing GPT-4o-mini for simpler questions to find a cost-efficient solution to the easy and hard tasks. Even if GPT-4o-mini has only a marginal chance of adequately addressing hard questions, it may still be cost-effective to initially attempt these prompts using GPT-4o-mini and resort to o3-mini only when necessary, due to the substantial price difference between the two models.
The example above highlights the necessity of incorporating both pricing considerations and the possibility of multiple prompt assignments into the online prompt-to-model assignment process. To the best of our knowledge, current model assignment methods do not account for pricing disparities between LLMs and generative AI services, which can be significant even among models offered by the same provider which may exhibit price differences as significant as $250$ times (e.g., OpenAI’s GPT4.5 at \$$150$ PMT vs. GPT4o-mini at \$$0.6$ PMT).
To address this gap, we introduce PromptWise, a novel cost-aware prompt-to-model assignment method. PromptWise is designed as an online learning algorithm that explicitly accounts for two key factors: (1) the prompt-based pricing and performance metrics of the models, and (2) the option for multiple assignments of a single prompt, provided the total cost remains economically justified based on previous observations. More concretely, PromptWise functions as a budget-aware contextual bandit algorithm, allowing for dynamic multiple assignments within an online learning setting.
At its core, PromptWise leverages the established Upper Confidence Bound (UCB) framework (Auer, 2000; Bubeck and Cesa-Bianchi, 2012). Unlike standard contextual bandit regret formulations (Langford and Zhang, 2007; Li et al., 2010; Chu et al., 2011), which simply sum the losses across iterations, PromptWise adopts a unique sum-min regret structure. Here, each summation term represents the minimal performance regret across multiple assignments of a single prompt, while the cumulative cost penalty is aggregated over all iterations. This formulation ensures that only the best achievable performance for each prompt contributes to the performance regret, whereas costs accumulate in the conventional manner.
We empirically validate the PromptWise algorithm on both simulated and real-world tasks involving prompt-guided image generation and various LLM services. Simulated scenarios, grounded by known optimal assignments, demonstrate PromptWise’s rapid adaptability in cost-effectively addressing prompts. The algorithm exhibits human-like strategic behavior by first making attempts on cheaper models, alternating to more expensive options only if the lower-cost models fail to adequately address the prompt. Real-world experiments encompass Sudoku and chess puzzle-solving with varying complexity, as well as code generation and translation tasks from various benchmarks, including HumanEval (Chen et al., 2021), HumanEval-X (Zheng et al., 2023), and BigCodeBench (Zhuo et al., 2024). Our results confirm PromptWise’s ability to significantly outperform existing cost-unaware selection methods, achieving superior budget-performance trade-offs. Notably, compared to always selecting the highest-cost model, PromptWise reduces total expenses substantially while maintaining or even enhancing overall prompt-handling effectiveness. In summary, our main contributions are as follows:
  • Highlighting the role of cost considerations and multiple prompt assignments in sequential prompt tasks.
  • Introducing PromptWise, a cost-aware contextual bandit framework optimized for balancing cost and performance with the possibility of multiple assignments of every prompt.
  • Defining a novel sum-min regret metric, which trades off between the maximum obtained reward and the cumulative cost per step.
  • Demonstrating PromptWise’s practical efficacy through several real-world and synthetic experiments, validating its effectiveness in cost-aware selection.
2 Related Work
Generalized linear contextual bandits. Contextual bandit (CB) is a popular framework for sequential decision-making with side information, which has been widely used in recommendation systems and advertising. A relevant assumption in the CB setup is that the expected reward conditioned to an incoming context admits a generalized linear model (GLM) (Filippi et al., 2010; Li et al., 2017). Particularly, the logistic bandit framework (Faury et al., 2020; Dong et al., 2019) has demonstrated stronger performance in applications where rewards are binary.
Bandit problems with cost constraints. A line of study considers multi-armed bandits (MAB) in a cost-aware setup, where each arm is associated with a cost distribution. The learner aims to not only optimize the cumulative return but also minimize the cost of pulling the arms. Ding et al. (Ding et al., 2013) study cost-aware bandits with a constraint on the total paying cost. In Multi-fidelity MAB (Kandasamy et al., 2016; Wang et al., 2023), the learner may choose to observe (pseudo-)rewards of an arm at some fidelity level and incur the corresponding cost. Recently, Kanarios et al. (Kanarios et al., 2024) study cost-aware best-arm identification, where the goal is to identify the largest reward arm using the minimum expected cost. To the best of our knowledge, cost-aware bandit algorithms in the literature consider optimizing a linear utility, e.g., the cumulative return minus (a regularized) cumulative cost.
Online learning methods for training and selecting generative models. A line of research incorporates online learning and reinforcement learning in the training and selecting generative models. Grnarova et al. (Grnarova et al., 2017) formulate the problem of training a Generative Adversarial Network (GAN) as finding a mixed strategy in a zero-sum game. They propose an online learning-based training algorithm, which provably converges to an equilibrium. Daskalakis et al. (Daskalakis et al., 2018) show that optimistic mirror descent  (OMD) can address the limit cycling issue in training GANs. More recent work proposes the reinforcement learning from human feedback (RLHF) pipeline for training large language models (LLMs) (Ouyang et al., 2022; Rafailov et al., 2023; Munos et al., 2024; Dong et al., 2024) and image-based generative models (Yuan et al., 2023; Xu et al., 2023). Regarding the selection of generative models, recent work proposes online learning-based algorithms (Hu et al., 2025) to a optimize performance score among a group of generative models.
3 Cost-Aware Contextual MAB
Interaction protocol. In this section, we introduce the framework of Cost-Aware Contextual MAB (CA-CMAB). The interaction protocol is given in Protocol 11. At each step $t\in[T]$, where $[T]:=\{1,2,\cdots,T\}$, a context $x_{t}\in{\mathcal{X}}\subseteq{\mathbb{B}}^{d}:=\{x\in{\mathbb{R}}^{d}:\|x\|_{2}\leq 1\}$ (e.g., the prompt input from the user) is drawn from some contextual distribution $p_{0}$. Different from the standard contextual bandit setting (Langford and Zhang, 2007), the learner may interact with context $x_{t}$ for multiple rounds by taking a sequence of actions in $\bar{{\mathcal{A}}}:=\{a_{0}\}\cup{\mathcal{A}}$, where $a_{0}$ is the null action and ${\mathcal{A}}$ is the set of arms. Specifically, after pulling an arm $a_{t,i}\in{\mathcal{A}}$ at the $i$-th round, the learner observes a reward $r_{t,i}\sim R_{a_{t,i}}(x_{t})$ and incurs a fixed cost $c_{t,i}:=c_{a_{t,i}}$. The environment moves on to the next step either when the learner takes the null action $a_{0}$ or the round budget is used up (i.e., $i=\tau_{\max}$). We denote by $\tau_{t}\in[\tau_{\max}]:=\{0,1,\cdots,\tau_{\max}\}$ the (random) number of arms pulled within step $t$. For convenience, we define $a_{t,\tau_{t}+1}=:a_{0}$. In addition, we assume that taking the null action yields a zero reward and cost, i.e., $R_{a_{0}}=c_{a_{0}}=1$. In this paper, we focus on the scenario where the rewards are binary.
Assumption 1 (Binary reward).
For any arm $a\in{\mathcal{A}}$, the reward distribution follows $\text{Ber}(q_{a}(x))$ for any incoming context $x\in{\mathcal{X}}$, where the success probability $q_{a}(x)\in[0,1]$.
Remark 1.
The assumption of binary reward has been widely adopted in the bandit literature, including Bernoulli bandits (Berry, 1972; Agrawal and Goyal, 2012) and the more recent logistic bandits (Dong et al., 2019; Faury et al., 2020). In practice, a binary reward suits applications such as puzzle-solving (i.e., correct or wrong) and code generation (i.e., pass or not pass).
1:step $T$, context distribution $p_{0}$, set of actions $\bar{{\mathcal{A}}}=\{a_{0}\}\cup{\mathcal{A}}$, cost $\{c_{a}\}_{a\in\bar{{\mathcal{A}}}}$, reward distributions $\{R_{a}\}_{a\in\bar{{\mathcal{A}}}}$, cost parameter $\lambda>0$, round budget $\tau_{\max}\in{\mathbb{N}}_{+}$.
2:for step $t=1,2,\cdots,T$ do
3:     The environment draws context $x_{t}\sim p_{0}$.
4:     Play action $a_{t,1}$ and expend $c_{a_{t,1}}$.
5:     Set counter $\tau_{t}\leftarrow 0$.
6:     while not playing $a_{0}$ and $\tau_{t}\leq\tau_{\max}$ do
7:         Play action $a_{t,\tau_{t}+1}$ and expend $c_{a_{t,\tau_{t}+1}}$.
8:         The environment draws reward $r_{t,\tau_{t}}\sim R_{a_{t,\tau_{t}}}(x_{t})$.
9:         Set $\tau_{t}\leftarrow\tau_{t}+1$ if $a_{t,\tau_{t}+1}\neq a_{0}$.
10:     end while
11:end for
Protocol 1 Cost-Aware Contextual MAB (CA-CMAB)
Policy and cost-aware utility. The learner interacts with the environment for $T$ steps. Upon receiving context $x_{t}\in{\mathcal{X}}$ at step $t$, the learner plays (a sequence of) actions according to some policy $\pi_{t}$, which may depend on the previous observations. The learner aims to find a policy that maximizes the following utility function, which trades off between the maximum obtained reward and the cumulative paid cost:
$\pi^{\star}(x_{t}):=\operatorname*{arg\,max}_{\pi}\left\{u^{\pi}(x_{t}):=\mathbb{E}_{\pi}\left[\max_{i=1,\cdots,\tau_{t}}r_{t,i}-\lambda\sum_{i=1}^{\tau_{t}}c_{t,i}\right]\right\},$ (1)
where $\lambda>0$ is the cost parameter.
Remark 2.
We highlight in Figure 11 the differences and unique challenges of the proposed CA-CMAB framework compared to the standard contextual bandit (CB) settings. CA-CMAB allows the learner to interact with a context multiple times before moving on to the next step. To optimize the performance, the learner faces a dilemma between making additional attempts to obtain a higher reward and paying extra cost for pulling the arms. On the other hand, the utility (1) is often a non-linear function for round budget $\tau_{\max}>1$. Hence, the optimal policy can be non-trivially different from the one in a MAB or CB environment (i.e., $\tau_{\max}=1$), where the learner can simply play the arm that maximizes $\mathbb{E}_{R_{a}(x)}[r]-\lambda c_{a}$.
Under the assumption of binary rewards and unlimited round budget, we show that there exists a deterministic optimal policy. The proof can be found in Appendix A.1.
Proposition 1 (Oracle).
Under Assumption 1 and with unlimited round budget, an optimal policy (1) is as follows: for any incoming context $x\in{\mathcal{X}}$, it takes action $a^{\star}(x)$ if a reward of $1$ is not observed, which is given by
$a^{\star}(x)=\begin{cases}a_{0},&\textup{if }\max_{a\in{\mathcal{A}}}\{q_{a}(x)-\lambda c_{a}\}\leq 0,\\ \operatorname*{arg\,min}_{a\in{\mathcal{A}}}\frac{c_{a}}{q_{a}(x)},&\textup{o.w.},\end{cases}$ (2)
where $q_{a}(x)$ and $c_{a}$ are the success probability conditioned to context $x$ and the (fixed) cost for any arm $a\in{\mathcal{A}}$, respectively. The expected total cost of this policy is $\frac{c_{a^{\star}(x)}}{q_{a^{\star}(x)}}$, and the expected utility is given by
$u^{\star}(x)=\mathbbm{1}\left[a^{\star}(x)\neq a_{0}\right]\cdot\left(1-\lambda\cdot\frac{c_{a^{\star}(x)}}{q_{a^{\star}(x)}}\right).$ (3)
Moreover, to handle the relationship between a (continuous) context $x\in{\mathcal{X}}$ and the conditional reward $R_{a}(x)$, we work under the following assumption.
Assumption 2 (Disjoint logistic-linear reward).
For any arm $a\in{\mathcal{A}}$, there exists $\theta_{a}^{\star}\in{\mathbb{R}}^{d}$ such that the success probability $q_{a}(x)=\mu(\langle\theta_{a}^{\star},x\rangle)$ for any context $x\in{\mathcal{X}}\subseteq{\mathbb{R}}^{d}$, where $\mu(z):=(1+\exp(-z))^{-1},z\in{\mathbb{R}}$ is the sigmoid function.
Figure 2: Interaction protocol of the proposed CA-CMAB framework (Below) and standard contextual bandits (Above): CA-CMAB allows multiple model assignments to adequately address the input prompt, while at the cost of paying extra service fees.
4 The PromptWise Algorithm
In this section, we propose PromptWise as an online learning-based algorithm to address the CA-CMAB framework. For an input context, PromptWise performs cost-aware optimization by recursively computing the oracle action (2) on a set of estimated success probabilities. As these estimates are calculated from the previous observations and may deviate from their true values, directly plugging them into Equation (2) can result in a sub-optimal action. To resolve this issue, PromptWise leverages the Upper Confidence Bound (UCB) framework (Auer, 2000; Bubeck and Cesa-Bianchi, 2012) in the online learning literature to address the estimation error.
Overview of PromptWise. We are ready to introduce PromptWise (presented in Algorithm 17). In the exploration phase (Lines 1-4), PromptWise plays each arm for $\tau_{\exp}$ steps, where $\tau_{\exp}\in{\mathbb{N}}_{+}$ is a tuning parameter. Upon receiving the rewards, PromptWise calculates the maximum-likelihood estimate (MLE) of the weight vectors $\{\theta_{a}^{\star}:a\in{\mathcal{A}}\}$, which are used to estimate the success probabilities (Lines 3 and 6):
$\operatorname*{arg\,max}_{\theta\in{\mathbb{R}}^{d}}\widehat{\ell}_{a}(\theta):=\sum_{(X,R)\in{\mathcal{D}}_{a}}\left(R\cdot\log\mu(\langle\theta,X\rangle)+(1-R)\cdot\log(1-\mu(\langle\theta,X\rangle))\right),$ (4)
which simplifies to solving Equation (5). (The detailed derivation can be found in Appendix B.1.) At any subsequent step (Lines 5-16), PromptWise computes the oracle action (2) on a set of estimated success probabilities and interacts with the environment (Lines 8-11). Particularly, these estimates are optimistic such that $\widehat{q}_{a}$ computed in Equation (6) upper bounds the groundtruth value $q_{a}(x_{t})$ with high probability. After receiving a new observation, PromptWise performs inner-step updates to both the MLE and the estimated success probability for the played arm (Line 13). Hence, PromptWise can leverage observations in the early rounds to adjust arm selections in later rounds.
1:step $T$, cost parameter $\lambda>0$, regression dataset ${\mathcal{D}}_{a}\leftarrow\varnothing$ for $a\in{\mathcal{A}}$, $V_{a}\leftarrow\bm{0}_{d\times d}$, round budget $\tau_{\max}$, tuning parameter $\tau_{\textup{exp}}$ and $\alpha$
2:for $a\in{\mathcal{A}}$ do
3:     Collect ${\mathcal{D}}_{a}\leftarrow{\mathcal{D}}_{a}\cup\{(x_{s},r_{s})\}_{s=1}^{\tau_{\exp}}$ by pulling arm $a$ (and moving on to the next context) upon observing contexts $\{x_{s}\}_{s=1}^{\tau_{\exp}}$ drawn from the environment.
4:     Calculate the maximum-likelihood estimate (MLE) $\widehat{\theta}_{a}$ by solving
$\sum_{(X,R)\in{\mathcal{D}}_{a}}(R-\mu(X^{\top}\theta))X=0.$ (5)
5:end for
6:for $t=|{\mathcal{A}}|\tau_{\textup{exp}}+1,\cdots,T$ do
7:     Observe context $x_{t}$ and compute the estimated success probability for each arm $a\in{\mathcal{A}}$:
$\widehat{q}_{a}\leftarrow\mu\left(x_{t}^{\top}\widehat{\theta}_{a}+\alpha\|x_{t}\|_{V_{a}^{-1}}\right),\textup{ where }V_{a}\leftarrow\sum_{(X,R)\in{\mathcal{D}}_{a}}XX^{\top}.$ (6)
8:     while not playing null action $a_{0}$ do
9:         if $\max_{a\in{\mathcal{A}}}\{\widehat{q}_{a}-\lambda\cdot c_{a}\}\leq 0$ or hitting the round budget or receiving a reward of 1 then
10:              Play null action $\widehat{a}\leftarrow a_{0}$ and move on to the next step.
11:         else
12:              Play arm $\widehat{a}\leftarrow\operatorname*{arg\,min}\frac{c_{a}}{\widehat{q}_{a}}$ and receive reward $r$.
13:              Update the regression dataset ${\mathcal{D}}_{\widehat{a}}\leftarrow{\mathcal{D}}_{\widehat{a}}\cup\{(x_{t},r)\}$.
14:              Recompute the MLE $\widehat{\theta}_{a}$ and the estimate $\widehat{q}_{\widehat{a}}$ by Equations (5) and (6), respectively.
15:         end if
16:     end while
17:end for
Algorithm 1 PromptWise
Remark 3.
To provide theoretical justification to the proposed approach, we show that a variant of PromptWise (presented in Algorithm 15 in the Appendix), which performs MLE updates only at the end of a step, can attain a sublinear regret. Details can be found in Theorem 8 in the Appendix. On the other hand, PromptWise in Algorithm 17 recursively updates the estimation in a round-by-round manner, which enables effective adaptation to an unobserved context. In Appendix C.1, we compare these two algorithms on all the setups in the Numerical Experiments Section 5. The results show that inner-step updates are key to the superior performance of the PromptWise algorithm.
5 Numerical Experiment
We report numerical results of the proposed PromptWise algorithm on the selection among a group of large language models (LLM). In the experiments, we test PromptWise (Algorithm 17) and the baseline methods on various tasks, including solving random Sudoku games, code completion, and code translation. Details and additional results can be found in Appendix C.
Summary. We highlight main findings in our numerical experiments. First, the proposed PromptWise algorithm can effectively balance between cost and model performance conditioned to the input prompt. Specifically, PromptWise strategically queries answers from lower-cost models first, proceeding to more expensive options only if the cheaper models cannot adequately address the prompt. Second, compared to the baseline methods, PromptWise can usually attain higher success rates (in addressing the input prompt) and/suffering lower cost, which showcases its superiority in cost-performance optimization. Third, PromptWise can adapt to new models and prompts in the selection process, which demonstrates its potential and practical efficacy in real-world applications.
5.1 Experimental Setup
Baselines. We compare the proposed PromptWise algorithm (Algorithm 17) to the following baseline methods: 1) Random-till-Succeed (RtS): randomly selecting a model until a satisfactory answer is obtained or the round budget is used up, 2) Greedy-till-Succeed (GtS): selecting the model that balances between the (average) success rate and the cost, i.e., $\operatorname*{arg\,max}_{a\in{\mathcal{A}}}\{\widehat{q}_{a}-\lambda c_{a}\}$, until a satisfactory answer is obtained or the budget is used up, 3) we also include PAK-UCB(-poly3) (Hu et al., 2025) as a baseline, which is a kernel-based contextual bandit algorithm that selects a model according to the input prompt to maximize the expected reward. In addition, we report results for four additional baselines: 4) Random (’- -’): randomly selecting a model and then moving on to the next prompt, 5) Greedy (’- -’): selecting the model that balances between the (average) success rate and the cost and then moving on to the next prompt, 6) Lowest-cost: always selecting the cheapest model, and 7) Highest-cost: always selecting the most expensive model. In the experiments, all the algorithms explore each arm at least once, except for the Lowest-cost and Highest-cost baselines.
Performance metrics. We evaluate the algorithmic performance using the following metrics: 1) Avg.Utility: utility per step attained by the algorithm, i.e., $T^{-1}\sum_{t=1}^{T}u(x_{t})$, 2) Avg.Cost: cost per step paid by the algorithm, i.e., $T^{-1}\sum_{t=1}^{T}\sum_{i=1}^{\tau_{t}}c_{t,i}$, 3) Approx.OPR: ratio that the algorithm picks the optimal action conditioned to the prompt, i.e., $T^{-1}\sum_{t=1}^{T}(\tau_{t}^{-1}\sum_{i=1}^{\tau_{t}}\mathbbm{[}a_{t,i}=a^{\star}(x_{t})]+\mathbbm{1}[a_{t,1}=a_{0}]\cdot\mathbbm{1}[a^{\star}(x_{t})=a_{0}])$. In addition, we report results of 3) Avg.Success: the proportion of steps where the selected model successfully addresses the prompt.
Candidate models and costs. For random Sudoku games (Task 1), we consider the following models: GPT-4o, GPT-4o-mini, o3-mini, Deepseek-Chat, and Qwen-Plus. For code generation tasks (Task 2&3), the candidate models include GPT-4o-mini, Claude-3.5-Haiku, Gemini-2.5-Flash-preview, Deepseek-Chat, and Qwen-Plus. The model costs are calculated as the total price of 1M input and output tokens (See Table 2 in the Appendix).
Hyperparameters. In the experiments, we set hyperparameters of PromptWise as $\tau_{\exp}=1$ and $\alpha=\log(|{\mathcal{A}}|/\delta)$, where $\delta=0.05$ is the failure probability. The cost parameter and round budget are set to be $\lambda=0.05$ and $\tau_{\max}=5$.
5.2 Numerical Results
Task 1: Random Sudoku games. In this experiment, we utilize LLMs to solve randomly generated $9\times 9$ Sudoku games. The LLM is given the initial board with $5$-$30$ blank entries. The sample prompt and answers are visualized in Figures 13 and 16 in the Appendix. For the PromptWise and PAK-UCB algorithms, we utilize a $5$-dimensional one-hot embedding as the input context, which is based on the number of blank entries in the initial board. Our results show that PromptWise can significantly outperform the baseline methods and achieve a good balance between the utility and the cost (Figures 3 and 19). To further illustrate this idea, we detail the model selection of PromptWise conditioned to the problem difficulty. In Table 1, we present the average $\textup{pass}@1:=\frac{c}{n}$ rate conditioned to three difficulty levels, where we sample $n=10$ solutions from each model for each Sudoku task and $c$ is the number of correct solutions among these ten trials. Interestingly, PromptWise queries cheaper models with slightly lower success rate (e.g., GPT-4o-mini and Deepseek-Chat) for easier Sudoku tasks while alternating to the expensive o3-mini model for harder Sudoku tasks.
Model Price Easy (5-15 blanks) Medium (15-25 blanks) Hard (25-30 blanks)
GPT-4o-mini $0.75$ $75.20\%$ $23.46\%$ $6.81\%$
[…]
Table 1: Pass@1 rate and model selection by PromptWise on random Sudoku games (Task 1). For each problem instance, we prompt an LLM for ten trials to compute the pass@1 rates. Model with the best performance in marked in bold text, while the selection by PromptWise is highlighted in orange.
Figure 3: Random Sudoku games (Task 1): GPT-4o, GPT-4o-mini, o3-mini, Deepseek-Chat, and Qwen-Plus. Results are averaged over $20$ trials.
Task 2: Code completion. In these experiments, we select LLMs to complete Python code. We consider tasks from two benchmarks: HumanEval (Chen et al., 2021) and BigCodeBench (Zhuo et al., 2024). Specifically, the model is provided with a function description (written in natural language) and is asked to complete the code following the function declaration (written in Python). Sample prompt and answers are visualized in Figures 14 and 17 in the Appendix. We utilize the CodeBERT embedding as the input context. More details can be found in Part 3 of Appendix C.2. Our results show that PromptWise can consistently outperform the baseline methods across the benchmarks (Figures 420 and 21).
Figure 4: Code completion on BigCodeBench (Task 2): GPT-4o-mini, Claude-3.5-Haiku, Gemini-2.5-Flash-preview, Deepseek-Chat, and Qwen-Plus. We utilize the first $164$ tasks to align with the HumanEval and HumanEval-X benchmark. Results are averaged over $20$ trials.
Task 3: Code translation. In this task, we prompt LLMs to translate codes between different programming languages. Specifically, the LLM is provided with a function written in the source programming language (e.g., C++) without any further description/explanation. Then, the model is asked to translate the code following the function declaration in the target programming language (e.g., Java). Sample prompt and answers are visualized in Figures 15 and 18. The tasks are randomly drawn from the HumanEval-X benchmark (Zheng et al., 2023).
Figure 5: Code translation on the HumanEval-X benchmark (Task 3): GPT-4o-mini, Gemini-2.5-Flash-preview, Claude-3.5-Haiku, Deepseek-Chat, and Qwen-Plus. Results are averaged over $20$ trials.
Performance synthetic text-to-image generation. We test the performance of PromptWise on text-to-image (T2I) models on a synthetic setup, where prompts are randomly drawn from five categories in the MS-COCO dataset (Lin et al., 2015). The five arms synthesized from Stable Diffusion v2 (Rombach et al., 2022) are visualized in Figure 23 in the Appendix, where the $i$-th arm can output high-quality images for the first $i$ prompt types. For each input prompt, the learner aims to generate an image up to some satisfactory level (i.e., the CLIPScore is higher than $30$). Results are summarized in Figure 24 in the Appendix.
Adaptation to new models and prompts. We also test PromptWise on the code translation task under a scenario where new models and prompts are introduced. This scenario is common in practice due to version updates and the release of new models. On the other hand, the interests of a typical user may vary from time to time. In the first setup (Figures 6 and 25), the model is asked to translate between C++ and Java code. At the beginning, the learner can access two models, i.e., GPT-4o-mini and Gemini-2.5-Flash-preview. In the process, Claude-3.5-Haiku, Deepseek-Chat, and Qwen-Plus are added to the pool after each $200$ steps. In the second setup (Figure 7 and 26), new translation tasks are introduced in the model selection process. Specifically, the task pool initially contains Java-to-C++ translation problems. Later on, C++-to-Java and Python-to-C++ translation tasks are introduced. Our results show that the proposed PromptWise can adapt to new models and prompts in the model selection process and significantly outperform the baselines.
Figure 6: Adaptation to new models on code translation task: The model pool contains two LLMs initially, i.e., GPT-4o-mini and Gemini-2.5-Flash-preview. In the process, Claude-3.5-Haiku, Deepseek-Chat, and Qwen-Plus are added to the pool after each $200$ steps. Results are averaged over 20 trials.
Figure 7: Adaptation to new prompts on code translation task: GPT-4o-mini, Claude-3.5-Haiku, Gemini-2.5-Flash-preview, Deepseek-Chat, and Qwen-Plus. Initially, the LLM is asked to solve a C++ to Java translation. Then, Java-to-C++ and Python-to-C++ tasks are introduced after each $250$ steps. Results are averaged over 20 trials.
6 Conclusion
In this work, we focused on the task of prompt assignment to generative models in a cost-aware online scenario, where the goal is to optimize a utility function that trades off between the best attained performance and the cumulative cost of model queries. We proposed PromptWise, an online cost-aware prompt-to-model assignment method. We demonstrated PromptWise’s practical results on both simulated and real-world tasks involving prompt-guided image generation and various LLM services, including Sudoku-solving and code generation tasks. In our experiments, PromptWise could significantly improve utility over existing cost-unaware selection methods, achieving a better budget-performance trade-off. We note that our numerical experimentation is constrained to coding and puzzle-solving tasks where the LLM’s output can be evaluated using standard metrics. Extending the applications of cost-aware online selection of generative models to more general tasks for LLMs and text-to-video generation models will be a relevant future direction. In addition, combining the online selection process in PromptWise and other cost-aware bandit algorithms with the offline model selection methods to further improve the cost-performance trade-off will be another interesting direction for future studies.
Appendix A Proofs in Section 3
A.1 Oracle with Unlimited Budget
Proof.
Fix a context $x$ and rewrite $q_{a}:=q_{a}(x)$ for convenience. The problem can be formulated as an infinite-horizon Markov Decision Process (MDP) with three states ${\mathcal{S}}=\{0,1,{\mathcal{T}}\}$, where 00 and $1$ represent $\max\{r_{t,1},\cdots,r_{t,i}\}$ up to the $i$-th round, and ${\mathcal{T}}$ is the terminal state. The transition kernel is given by
$P(s^{\prime}|s,a)=\begin{cases}1,&\text{if }(s,a,s^{\prime})\in\{(s,a_{0},{\mathcal{T}})\}_{s\in{\mathcal{S}}}\cup\{(1,a,1)\}_{a\in{\mathcal{A}}}\\ q_{a},&\text{o.w.}\end{cases},$
and the reward function is defined as $r(s,a,s^{\prime})=u(s^{\prime})-u(s)-\lambda c_{a}$, where $u(0)=u({\mathcal{T}})=0$ and $u(1)=1$. The agent aims to maximize the cumulative reward $\sum_{i=1}^{\infty}r(s_{i},a_{i},s^{\prime}_{i})$. It is easy to verify that the optimal value function $V^{\star}({\mathcal{T}})=V^{\star}(1)=0$. At state 00, it holds that $Q^{\star}(0,a)=q_{a}(1-\lambda c_{a})+(1-q_{a})(-\lambda c_{a}+V^{\star}(0))$ and $V^{\star}(0)=\max_{a\in\bar{{\mathcal{A}}}}Q^{\star}(0,a)$. Hence, we have
$V^{\star}(0)=\max_{a\in\bar{{\mathcal{A}}}}\left\{q_{a}-\lambda c_{a}+(1-q_{a})V^{\star}(0)\right\},$ (7)
and we can derive $V^{\star}(0)=\max\{0,\frac{q_{a^{\star}}-\lambda c_{a^{\star}}}{q_{a^{\star}}}\}$, where
$a^{\star}=\operatorname*{arg\,max}_{a\in{\mathcal{A}}}\{q_{a}-\lambda c_{a}+(1-q_{a})V^{\star}(0)\}=\begin{cases}a_{0}&\text{if }V^{\star}(0)=0,\\ \operatorname*{arg\,min}_{a\in{\mathcal{A}}}\frac{c_{a}}{q_{a}}&\text{o.w.}\end{cases}$
We conclude the proof. ∎
Appendix B Proofs in Section 4
B.1 Maximum-Likelihood Estimate
Proof.
We show that the maximum-likelihood estimate (4) can be derived by solving Equation (5). Note that
^$\widehat{\ell}_{a}(\theta)=$ $\sum_{(X,R)\in{\mathcal{D}}_{a}}R\cdot\log\frac{\mu(\langle\theta,X\rangle)}{1-\mu(\langle\theta,X\rangle)}+\log(1-\mu(\langle\theta,X\rangle))$
$=$ $\sum_{(X,R)\in{\mathcal{D}}_{a}}R\cdot\langle\theta,X\rangle+\log(1-\mu(\langle\theta,X\rangle)).$
By taking the derivative w.r.t $\theta$, we derive
^$\widehat{\ell}^{\prime}_{a}(\theta)=\sum_{(X,R)}R\cdot X+\frac{-\mu^{\prime}(\langle\theta,X\rangle)}{1-\mu(\langle\theta,X\rangle)}X=\sum_{(X,R)}(R-\mu(\langle\theta,X\rangle))X,$
which concludes the proof. ∎
B.2 Sublinear Regret for PromptWise Variant
Definition of regret. In this section, we provide theoretical justification to the proposed PromptWise algorithm. Specifically, we show that a PromptWise variant attains sublinear regret, which is given by
$\textup{Regret}(T):=\mathbb{E}\left[\sum_{t=1}^{T}\left(u^{\star}(x_{t})-u^{\pi_{t}}(x_{t})\right)\right],$ (8)
where $\pi_{t}$ is the policy at step $t$, and we denote by $u^{\star}(x):=u^{\pi^{\star}(x)}(x)$ the optimal utility (1) conditioned to any context $x\in{\mathcal{X}}$. The regret metric is widely adopted in the online learning literature [Bubeck and Cesa-Bianchi, 2012]. Note that any algorithm that attains sublinear regret guarantees (asymptotic) convergence to the optimal policy, i.e., $\frac{1}{T}\mathbb{E}\left[\sum_{t=1}^{T}\left(u^{\star}(x_{t})-u^{\pi_{t}}(x_{t})\right)\right]\rightarrow 0$.
PromptWise variant. Next, we present the variant algorithm in Algorithm 15 below. Note that the only difference is that Algorithm 15 performs MLE updates only at the end of a step (Line 12), instead of doing it each after observing a reward (Lines 12-13 in Algorithm 17). As a result, Algorithm 15 either plays the null action $a_{0}$ upon observing an input context and moves on to the next step (Lines 7-8) or pulls an arm deterministically until succeeded or the round budget has been used up (Lines 10).
1:step $T$, cost parameter $\lambda>0$, regression dataset ${\mathcal{D}}_{a}\leftarrow\varnothing$ for $a\in{\mathcal{A}}$, $V_{a}\leftarrow\bm{0}_{d\times d}$, tuning parameter $\tau_{\textup{exp}}$ and $\alpha$
2:for $a\in{\mathcal{A}}$ do
3:     Collect ${\mathcal{D}}_{a}\leftarrow{\mathcal{D}}_{a}\cup\{(x_{s},r_{s})\}_{s=1}^{\tau_{\exp}}$ by pulling arm $a$ (and moving on to the next context) upon observing contexts $\{x_{s}\}_{s=1}^{\tau_{\exp}}$ drawn from the environment.
4:     Calculate the maximum-likelihood estimate (MLE) $\widehat{\theta}_{a}$ by solving
$\sum_{(X,R)\in{\mathcal{D}}_{a}}(R-\mu(X^{\top}\theta))X=0.$ (9)
5:end for
6:for $t=|{\mathcal{A}}|\tau_{\textup{exp}}+1,\cdots,T$ do
7:     Observe context $x_{t}$ and compute the estimated success probability for each arm $a\in{\mathcal{A}}$:
$\widehat{q}_{a}\leftarrow\mu\left(x_{t}^{\top}\widehat{\theta}_{a}+\alpha\|x_{t}\|_{V_{a}^{-1}}\right),\textup{ where }V_{a}\leftarrow\sum_{(X,R)\in{\mathcal{D}}_{a}}XX^{\top}.$ (10)
8:     if $\max_{a\in{\mathcal{A}}}\{\widehat{q}_{a}-\lambda\cdot c_{a}\}\leq 0$ then
9:         Play null action $\widehat{a}\leftarrow a_{0}$ and move on to the next step.
10:     else
11:         Play arm $\widehat{a}\leftarrow\operatorname*{arg\,min}_{a\in{\mathcal{A}}}\frac{c_{a}}{\widehat{q}_{a}}$ until receiving a reward of 1 or hitting the round budget.
12:         Update ${\mathcal{D}}_{\widehat{a}}\leftarrow{\mathcal{D}}_{\widehat{a}}\cup\{(x_{t},r_{i})\}_{i=1}^{\tau_{t}}$ where $r_{1},\cdots,r_{\tau_{t}}$ are the observed rewards.
13:         Recompute $\widehat{\theta}_{\widehat{a}}$ and $\widehat{q}_{\widehat{a}}$ for the played arm by Equations (9) and (10), respectively.
14:     end if
15:end for
Algorithm 2 PromptWise with Per-Step Updates
Regret analysis. To derive a regret bound for Algorithm 15, we state two assumptions for our analysis. The first one is a regularity assumption on the contextual distribution $p_{0}$, which is standard in the logistic bandit literature [Li et al., 2017]
Assumption 3.
There exists a constant $\sigma_{0}>0$ such that $\lambda_{\min}(\Sigma)\geq\sigma_{0}^{2}$, where $\Sigma:=\mathbb{E}_{X\sim p_{0}}[XX^{\top}]$.
Assumption 4 (Lower-bounded success rate).
There exists a constant $q_{0}$ such that $q_{a}(x)>q_{0}$ for any arm $a\in{\mathcal{A}}$ and context $x$ in the support of distribution $p_{0}$.
Let $\kappa:=\inf_{a\in{\mathcal{A}}}\{\mu^{\prime}(x^{\top}\theta_{a}^{\star}):\|x\|\leq 1,\|\theta-\theta_{a}^{\star}\|\leq 1\}$.
Theorem 1 (Regret).
Let the choice of $\tau_{\exp}$ follows from Lemma 1 with $B=1$ and a union bound over all the arms, i.e., $\tau_{\exp}=C\sigma_{0}^{-2}(d+\log(|{\mathcal{A}}|/\delta))$ for some universal constant $C>0$. Set $\alpha=\kappa^{-1}\sqrt{\frac{d}{2}\log(1+2\tau_{\max}T/d)+\log(|{\mathcal{A}}|/\delta)}$. Then, the regret of running Algorithm 15 with round budget $\tau_{\max}\geq\frac{\log(dq_{0}T^{-1/2})}{\log(1-q_{0})}$ for $T$ step is bounded by
$\textup{Regret}(T)\leq|{\mathcal{A}}|\tau_{\exp}+\widetilde{O}\left((\kappa^{-1}+1)d\sqrt{|{\mathcal{A}}|T}\right),$
where logarithmic factors are hidden in the notation $\widetilde{O}(\cdot)$.
Proof.
Optimism. Let $\widehat{q}^{t}_{a}:=\widehat{q}^{t}(x_{t})$ be the optimistic estimate (6) for any arm $a\in{\mathcal{A}}$ at (the beginning of) the $t$-th step. Note that $\widehat{q}^{t}_{a}$ is identical across the rounds within step $t$. We first show that, with probability at least $1-2\delta$, it holds that $\widehat{q}_{a}^{t}\geq q_{a}^{t}$ at any iteration $t\geq|{\mathcal{A}}|\tau_{\exp}+1$, where we denote by $q_{a}^{t}:=q_{a}(x_{t})$ the groundtruth expected value. By Lemma 1 and the choice of $\tau_{\exp}$, we have $\lambda_{\min}(V_{a})\geq 1$ for any $a\in{\mathcal{A}}$ with probability at least $1-\delta$. Let $\widehat{\theta}_{a}^{t}$ denote the maximum-likelihood estimate at the $t$-th iteration solving from Equation (5). Note that
$\widehat{q}_{a}^{t}=\mu\left(\langle x_{t},\widehat{\theta}_{a}^{t}\rangle+\alpha\|x_{t}\|_{(V_{a}^{t})^{-1}}\right)=\mu\left(\langle x_{t},\theta_{a}^{\star}\rangle+\langle x_{t},\widehat{\theta}_{a}^{t}-\theta_{a}^{\star}\rangle+\alpha\|x_{t}\|_{(V_{a}^{t})^{-1}}\right).$ (9)
By Lemma 2 and a union bound over all the arms, with probability at least $1-\delta$, we have
$\langle x_{t},\widehat{\theta}_{a}^{t}-\theta_{a}^{\star}\rangle+\alpha\|x_{t}\|_{(V_{a}^{t})^{-1}}\geq\|x_{t}\|_{(V_{a}^{t})^{-1}}\left(-\|\widehat{\theta}_{a}^{t}-\theta_{a}^{\star}\|_{V_{a}^{t}}+\alpha\right)\geq 0,$
where the second inequality holds by the Cauchy-Schwarz inequality. Hence, it holds that $\widehat{q}_{a}^{t}\geq q_{a}^{t}$ with probability at least $1-2\delta$. Therefore, we have
$\widehat{u}_{t}:=\mathbbm{1}[\widehat{a}_{t}\neq a_{0}]\cdot\left(1-\operatorname*{arg\,min}_{a\in{\mathcal{A}}}\frac{\lambda c_{a}}{\widehat{q}_{a}^{t}}\right)\geq u^{\star}_{\infty}(x_{t})\geq u^{\star}(x_{t}),$ (10)
where we denote by $u^{\star}_{\infty}(x_{t})$ the optimal utility given unlimited round budget, which is given in Equation (3).
Regret decomposition. To derive a upper bound for the regret, note that any step such that $\widehat{a}_{t}=a_{0}$ does not incur regret. Hence, we derive
$\text{Regret}(T)$
$=$ $\underbrace{\sum_{a\in{\mathcal{A}}}\sum_{s=1}^{\tau_{\exp}}\left(u^{\star}(x_{s})-(q_{a}(x_{s})-c_{a})\right)}_{\text{Exploration (Lines 1-8)}}+\underbrace{\sum_{t=|{\mathcal{A}}|\tau_{\exp}+1}^{T}\left(u^{\star}(x_{t})-u_{\textup{trunc}}^{\widehat{a}_{t}}(x_{t})\right)}_{\text{Lines 9-18}}$
$\leq$ $|{\mathcal{A}}|\tau_{\exp}+\sum_{t=|{\mathcal{A}}|\tau_{\exp}+1}^{T}\mathbbm{1}[\widehat{a}_{t}\neq a_{0}]\cdot\left(\widehat{u}_{t}-u^{\widehat{a}_{t}}_{\infty}(x_{t})+u^{\widehat{a}_{t}}_{\infty}(x_{t})-u_{\textup{trunc}}^{\widehat{a}_{t}}(x_{t})\right)$ (Inequality (10))
$\leq$ $|{\mathcal{A}}|\tau_{\exp}+\sum_{t=|{\mathcal{A}}|\tau_{\exp}+1}^{T}\mathbbm{1}[\widehat{a}_{t}\neq a_{0}]\cdot\left(\left(\frac{\lambda c_{\widehat{a}_{t}}}{q_{\widehat{a}_{t}}^{t}}-\frac{\lambda c_{\widehat{a}_{t}}}{\widehat{q}_{\widehat{a}_{t}}^{t}}\right)+\left|\frac{(1-q_{\widehat{a}_{t}}^{t})^{\tau_{\max}}(q_{\widehat{a}_{t}}^{t}-\lambda c_{\widehat{a}_{t}})}{q_{\widehat{a}_{t}}^{t}}\right|\right)$ (Lemma 4)
$\leq$ $|{\mathcal{A}}|\tau_{\exp}+\sum_{t=|{\mathcal{A}}|\tau_{\exp}+1}^{T}\mathbbm{1}[\widehat{a}_{t}\neq a_{0}]\cdot\left(q_{0}^{-2}(\widehat{q}_{\widehat{a}_{t}}^{t}-q_{\widehat{a}_{t}}^{t})+\frac{(1-q_{\widehat{a}_{t}}^{t})^{\tau_{\max}}}{q_{\widehat{a}_{t}}^{t}}\right),$ (Assumption 4)
where in the first inequality, we denote by $u_{\textup{trunc}}^{\widehat{a}_{t}}(x_{t})$ the utility of the truncated policy in Lines 7-13 of Algorithm 15, that is, playing the null action $a_{0}$ if $\widehat{u}_{t}\leq 0$, or otherwise, keep playing arm $\widehat{a}_{t}$ until receiving a reward of 1 or hitting the round budget. In addition, we define $u^{\widehat{a}_{t}}_{\infty}(x_{t}):=1-\lambda c_{\widehat{a}_{t}}/q_{\widehat{a}_{t}}$ to be the utility with unlimited budget. Note that the gap $|u^{\widehat{a}_{t}}_{\infty}(x_{t})-u_{\textup{trunc}}^{\widehat{a}_{t}}(x_{t})|$ can be directly derived from Lemma 4. In the last inequality, we utilize the fact that $\lambda c_{\widehat{a}_{t}}<\widehat{q}^{t}_{\widehat{a}_{t}}\leq 1$.
Bound $\widehat{q}_{\widehat{a}_{t}}^{t}-q_{\widehat{a}_{t}}^{t}$. Note that $\mu^{\prime}(z)\leq\frac{1}{4}$. From Equation (9), we derive
$\widehat{q}_{a}(x_{t})-q_{a}(x_{t})\leq\frac{1}{4}\left(\langle x_{t},\widehat{\theta}_{a}^{t}-\theta_{a}^{\star}\rangle+\alpha\|x_{t}\|_{(V_{a}^{t})^{-1}}\right)\leq\frac{\alpha}{2}\|x_{t}\|_{(V_{a}^{t})^{-1}}.$
Putting everything together. We have
$\text{Regret}(T)$
$<$ $|{\mathcal{A}}|\tau_{\exp}+q_{0}^{-2}\sum_{t=|{\mathcal{A}}|\tau_{\exp}+1}^{T}\mathbbm{1}[\widehat{a}_{t}\neq a_{0}]\cdot(\widehat{q}_{\widehat{a}_{t}}^{t}-q_{\widehat{a}_{t}}^{t})+T(1-q_{0})^{\tau_{\max}}q_{0}^{-1}$
$=$ $|{\mathcal{A}}|\tau_{\exp}+T(1-q_{0})^{\tau_{\max}}q_{0}^{-1}+q_{0}^{-2}\alpha\sum_{t=|{\mathcal{A}}|\tau_{\exp}+1}^{T}\mathbbm{1}[\widehat{a}_{t}\neq a_{0}]\cdot\|x_{t}\|_{(V_{a}^{t})^{-1}}.$
Let $n_{a}^{T+1}$ denote the size of the regression data set ${\mathcal{D}}_{a}$ of any arm $a\in{\mathcal{A}}$ after $T$ steps. Further, by Lemma 2, we have
$\sum_{t=|{\mathcal{A}}|\tau_{\exp}+1}^{T}\mathbbm{1}[\widehat{a}_{t}\neq a_{0}]\cdot\|x_{t}\|_{(V_{a}^{t})^{-1}}\leq$ $\sum_{a\in{\mathcal{A}}}\sqrt{2n_{a}^{T+1}d\log\left(\frac{\tau_{\max}T}{d}\right)}$
$\leq$ $\sqrt{2|{\mathcal{A}}|\tau_{\max}Td\log\left(\frac{\tau_{\max}T}{d}\right)}.$
By the choice of $\tau_{\max}\geq\frac{\log(dq_{0}T^{-1/2})}{\log(1-q_{0})}$, the total regret is bounded by
$\text{Regret}(T)\leq|{\mathcal{A}}|\tau_{\exp}+O(d\sqrt{T})+\widetilde{O}(\kappa^{-1}d\sqrt{|{\mathcal{A}}|T}),$
which concludes the proof. ∎
B.3 Auxiliary Lemmas
Lemma 1 ([Li et al., 2017, Proposition 1]).
Define $V_{n}:=\sum_{s=1}^{n}X_{s}X_{s}^{\top}$ where $X_{s}\in{\mathbb{R}}^{d}$ is drawn i.i.d from distribution $p_{0}$ with support in the unit ball. Further, let $\Sigma=\mathbb{E}_{X\sim p_{0}}[XX^{\top}]$ be the second moment matrix, and $B$ and $\delta>0$ be two positive constants. Then, there exist positive, universal constants $C_{1}$ and $C_{2}$ such that $\lambda_{\min}(V_{n})\geq B$ with probability at least $1-\delta$, as long as
$|{\mathcal{D}}_{a}|\geq\left(\frac{C_{1}\sqrt{d}+C_{2}\sqrt{\log(1/\delta)}}{\lambda_{\min}(\Sigma)}\right)^{2}+\frac{2B}{\lambda_{\min}(\Sigma)}.$
The following lemma is an adaptation of [Li et al., 2017, Lemma 3].
Lemma 2.
Let $V_{a}^{t}$ be the empirical second moment matrix for a fixed arm $a\in{\mathcal{A}}$ up to step $t$, and we denote by $n_{a}^{t}$ the size of the regression dataset ${\mathcal{D}}_{a}$. At any iteration $t\geq|A|\tau_{\exp}$, we denote by $\widehat{\theta}_{a}^{t}$ the maximum-likelihood estimator corresponding to any arm $a\in{\mathcal{A}}$. Suppose $\lambda_{\min}(V_{a}^{|{\mathcal{A}}|\tau_{\exp}+1})\geq 1$. For any $\delta\in[T^{-1},1)$, it holds that
$\|\widehat{\theta}_{a}^{t}-\theta_{a}^{\star}\|_{V_{a}^{t}}\leq\kappa^{-1}\sqrt{\frac{d}{2}\log\left(1+\frac{2n_{a}^{t}}{d}\right)+\log\frac{1}{\delta}}$
for all $t\geq|A|\tau_{\exp}$ with probability at least $1-\delta$.
Lemma 3 ([Li et al., 2017, Lemma 2]).
Let $\{X_{n}\}_{n=1}^{\infty}$ be a sequence in ${\mathbb{R}}^{d}$ satisfying $\|X_{n}\|\leq 1$. Define $X_{0}:=\bm{0}$ and $V_{n}:=\sum_{s=0}^{n-1}X_{s}X_{s}^{\top}$. Suppose there is an integer $m$ such that $\lambda_{\min}(V_{m+1})\geq 1$, then for all $s>0$,
$\sum_{n=m+1}^{m+s}\|X_{n}\|_{V_{n}^{-1}}\leq\sqrt{2nd\log\left(\frac{s+m}{d}\right)}.$
Lemma 4 (Truncated utility).
Let $R\sim\text{Ber}(q)$ be a Bernoulli random variable and $c>0$ be a constant. Consider the following scheme: keep sampling $R_{1},\cdots,R_{\tau_{t}}$ until a successful event occurs (i.e., $R_{\tau_{t}}=1$) or hitting the budget $\tau_{\max}$ (i.e., $\tau_{t}=\tau_{\max}$). Then, we have
$\mathbb{E}\left[\max_{i=1,\cdots,\tau_{t}}R_{i}-\tau_{t}c\right]=1-\frac{\lambda c}{q}-\frac{(1-q)^{\tau_{\max}}(q-\lambda c)}{q}.$
Proof.
Note that
$\mathbb{E}\left[\max_{i=1,\cdots,\tau_{t}}R_{i}-\tau_{t}c\right]=$ $\sum_{n=1}^{\tau_{\max}}(1-q)^{n-1}q(1-n\lambda c)+(1-q)^{\tau_{\max}}(-\lambda\tau_{\max}c)$
$=$ $1-\lambda c\sum_{n=1}^{\tau_{\max}}n(1-q)^{n-1}q-(1-q)^{\tau_{\max}}(1+\lambda\tau_{\max}c)$
$=$ $1-\lambda c\frac{1-(1-q)^{\tau_{\max}}(1+\tau_{\max}q)}{q}-\frac{(1-q)^{\tau_{\max}}q(1+\lambda\tau_{\max}c)}{q}$
$=$ $1-\frac{\lambda c}{q}-\frac{(1-q)^{\tau_{\max}}(q-\lambda c)}{q},$
which concludes the proof. ∎
Proposition 2 (Optimism).
Let $\{\widehat{q}_{a}(x):a\in{\mathcal{A}}\}$ be a set of optimistic estimates such that $\widehat{q}_{a}(x)\geq q_{a}(x)$ for any action $a\in{\mathcal{A}}$ and context $x\in{\mathcal{X}}$. We denote by $\widehat{u}(x)$ the utility solved from Equation (3) on the optimistic estimates. Then, we have $\widehat{u}(x)\geq u^{\star}_{\infty}(x)\geq u^{\star}(x)$ for any $x\in{\mathcal{X}}$.
1import itertools
2from random import shuffle
3
4def task_func(numbers=list(range(1, 3))):
5"""
6Calculates the average of the sums of absolute differences between each pair of consecutive numbers for all permutations of a given list. Each permutation is shuffled before calculating the differences.
7
8Args:
9 numbers (list): A list of numbers. Default is numbers from 1 to 10.
10
11Returns:
12 float: The average of the sums of absolute differences for each shuffled permutation of the list.
13
14Requirements:
15 - itertools
16 - random.shuffle
17
18Example:
19 $>$ result = task_func([1, 2, 3])
20 $>$ isinstance(result, float)
21 True
22"""
Appendix C Experimental Details and Additional Results
C.1 PromptWise and Its Variant (Alg.15)
(a) Sudoku-Solving (Task 1)
(b) Code completion on HumanEval (Task 2)
(c) Code completion on BigCodeBench (Task 2)
(d) Code translation on HumanEval-X (Task 3)
(e) Synthetic text-to-image generation
(f) Adaptation to new models
(g) Adaptation to new prompts
Figure 8: Performance of PromptWise and its variant with per-step updates (Algorithm 15).
C.2 Ablation Study
1. Cost parameter $\lambda$.
(a) $\lambda=0.01$
(b) $\lambda=0.05$ (default)
(c) $\lambda=0.1$
Figure 9: Ablation study on cost parameter $\lambda$.
2. Round budget $\tau_{\max}$.
(a) $\tau_{\max}=1$
(b) $\tau_{\max}=5$ (default)
(c) $\tau_{\max}=10$
Figure 10: Ablation study on round budget $\tau_{\max}$.
3. Input context. We test the PromptWise algorithm for different embeddings on the code completion task. Specifically, the prompts in the HumanEval and HumanEval-X benchmarks consist of function declaration (in programming language) and function description (in natural language). The sample prompt for task ’BigCodeBench/0’ in the HumanEval-X benchmark is visualized in Listing LABEL:lst:task-example, where the text in green color corresponds to function description.
We consider three types of embeddings: 1) Codebert [Feng et al., 2020] embedding with $768$ dimensions of the complete prompt (default), 2) Roberta [Liu et al., 2019] embedding with $768$ dimensions of the complete prompt, and 3) a concatenation of Codebert embedding of function declaration and Roberta embedding of function description, which results in $1536$ dimensions.
Figure 11: Ablation study on embeddings: Results on the HumanEval benchmark.
Figure 12: Ablation study on embeddings: Results on BigCodeBench.
C.3 Experimental Setup
All the experiments can be run on one single GeForce RTX 4090 Graphics Card.
Model costs.
Model Input / 1M Tokens (\$) Output / 1M Tokens (\$) Set Price
gpt-4o-2024-08-06 $2.50$ $10.00$ $12.50$
gpt-4o-mini-2024-07-18 $0.15$ $0.60$ $0.75$
o3-mini-2025-01-31222https://platform.openai.com/docs/pricing $1.10$ $4.40$ $5.50$
Gemini-2.5-Flash-preview (Non-thinking)  [^3] https://ai.google.dev/gemini-api/docs/pricing $0.15$ $0.60$ $0.75$
Claude-3.5-Haiku $3.00$ $15.00$ $18.00$
Claude-3.7-Sonnet [^4] https://www.anthropic.com/pricing#api $0.80$ $4.00$ $4.80$
Deepseek-Chat [^5] https://api-docs.deepseek.com/quick_start/pricing/. We use the standard prices. $0.27$ $1.10$ $1.37$
Qwen-Plus-2025-01-25666https://www.alibabacloud.com/help/en/model-studio/models $0.40$ $1.20$ $1.60$
Table 2: Price of API calls
Sample prompts and answers. We provide examples of the input prompt in different tasks (Figures 13-15). Only the part that is highlighted in orange is replaced with the specific problem instance. In addition, we visualize sample answers from several LLMs in Figures 16-18.
Figure 13: Sample prompt and answer in solving random Sudoku games (Task 1): The initial board is given where the empty cells are marked with ’.’. The LLM is asked to provide the finished board.
Figure 14: Sample prompt and answer in code completion task (Task 2): The declaration and functionality are provided. The LLM is asked to complete the code.
Figure 15: Sample prompt and answer in code translation task (Task 3): A function written in source programming language (e.g., C++) is given. The LLM is asked to translate the code into target language (e.g., Java), where the declaration part is provided.
Figure 16: Sample answers from the Qwen-Plus model on random Sudoku games (Task 1).
Figure 17: Sample answers from the Gemini-2.5-Flash-preview model on code completion task (Task 2). The problem instance is ’HumanEval/83’ in the HumanEval benchmark [Chen et al., 2021].
Figure 18: Sample answers from the GPT-4o-mini model on C++ to Java translation task (Task 2). The prompt is not displayed for brevity.
C.4 Additional Results
We report additional numerical results for the Random (’- -’), Greedy (’- -’), Lowest-cost, and Highest-cost baselines. In addition, we report performance for the Avg.Success metric.
Figure 19: Random Sudoku games (Task 1): GPT-4o, GPT-4o-mini, o3-mini, Deepseek-Chat, and Qwen-Plus. Results are averaged over $20$ trials.
Figure 20: Code completion on the HumanEval benchmark (Task 2): GPT-4o-mini, Claude-3.5-Haiku, Gemini-2.5-Flash-preview, Deepseek-Chat, and Qwen-Plus.
Figure 21: Code completion on BigCodeBench (Task 2): GPT-4o-mini, Claude-3.5-Haiku, Gemini-2.5-Flash-preview, Deepseek-Chat, and Qwen-Plus. Results are averaged over $20$ trials.
Figure 22: Code translation on the HumanEval-X benchmark (Task 3): GPT-4o-mini, Gemini-2.5-Flash-preview, Claude-3.5-Haiku, Deepseek-Chat, and Qwen-Plus. Results are averaged over $20$ trials.
Prompt Type (MS-COCO) Expert 1 Expert 2 Expert 3 Expert 4 Expert 5
Cost $0.5$ $1.0$ $1.5$ $2.0$ $2.5$
carrot Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
dog Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
cake Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
car Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
bowl Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 23: Arms in synthetic T2I generation: Prompts are randomly drawn from five categories in the MS-COCO dataset [Lin et al., 2015], including ’carrot’, ’dog’, ’cake’, ’car’, and ’bowl’. Each row and column display the generated images from an arm/expert conditioned to the prompt type.The $i$-th expert is constructed such that such that it outputs a (clean) image from Stable Diffusion v2 for the first $i$ prompt types, while outputting a white-noise-perturbed image for the rest of the prompt types.
Figure 24: Synthetic text-to-image (T2I) generation: The synthetic arms are visualized in Figure 23, where $i$-th arm is an expert in generating images for prompts of the first $i$ types. The reward for an generated image $I$ is defined to be $\mathbbm{1}[\text{CLIPScore}(I)\geq 30]$. Results are averaged over 20 trials.
Figure 25: Adaptation to new models on code translation task: The model pool contains two LLMs initially, i.e., GPT-4o-mini and Gemini-2.5-Flash-preview. In the process, Claude-3.5-Haiku, Deepseek-Chat, and Qwen-Plus are added to the pool after each $200$ steps. Results are averaged over 20 trials.
Figure 26: Adaptation to new prompts on code translation task: GPT-4o-mini, Claude-3.5-Haiku, Gemini-2.5-Flash-preview, Deepseek-Chat, and Qwen-Plus. Initially, the LLM is asked to solve C++-to-Java translation. Then, Java-to-C++ and Python-to-C++ tasks are introduced after each $250$ steps. Results are averaged over 20 trials.
1"""
2You are a very strong chess engine. Given the FEN r5k1/Rp3p1p/2b2qp1/3pr3/8/4P2P/2PN1PP1/Q3K2R b K - 0 19’, it is blacks turn to move. Provide the single best *legal* move in Standard Algebraic Notation (SAN). Examples: Nf3’, O-O’, Rxe5’, b8=Q’. Do not include commentary, checks (’+’), or checkmates (’#’). Output only the move."
3"""
C.5 Extension to Kernel-Based Methods
In this part, we discuss the extension of PromptWise to kernel-based methods. For an incoming context $x_{t}$ at the $t$-th step, we estimate the success probability $q_{a}(x_{t})$ for any arm $a\in{\mathcal{A}}$ by
$q_{a}(x_{t}|{\mathcal{D}}_{a})=\mu\left(\sum_{(X,R)\in{\mathcal{D}}_{a}}\beta_{a}^{X}K(X,x_{t})+{\mathcal{B}}\right),$
where $\beta_{a}^{X}\in{\mathbb{R}}$ is the weight corresponding to a datapoint $X\in{\mathcal{D}}_{a}$, $K:{\mathcal{X}}\times{\mathcal{X}}\to{\mathbb{R}}$ is the kernel function, and ${\mathcal{B}}:={\mathcal{B}}(x_{t}|{\mathcal{D}}_{a})$ is a bonus function. Specifically, the weight $\bm{\beta}\in{\mathbb{R}}^{|{\mathcal{D}}_{a}|}$ is solved by minimizing the regularized negative log-likelihood:
${\mathcal{L}}(\bm{\beta}_{a}):=-\sum_{(X,R)\in{\mathcal{D}}_{a}}\left(R\log q_{a}(X|{\mathcal{D}}_{a})+(1-R)\log(1-q_{a}(X|{\mathcal{D}}_{a}))\right)+\frac{\alpha}{2}\bm{\beta}_{a}^{\top}\bm{K}\bm{\beta}_{a}^{\top},$
where $\alpha>0$ is a tuning parameter that controls regularization and $\bm{K}:=\bm{K}({\mathcal{D}}_{a})\in{\mathbb{R}}^{|{\mathcal{D}}_{a}|\times|{\mathcal{D}}_{a}|}$ is the kernel matrix computed on the regression dataset. Moreover, the bonus function ${\mathcal{B}}:={\mathcal{B}}({\mathcal{D}}_{a})$ is added to ensure optimism for the estimated weight. In the implementation, we follow the work of [Hu et al., 2025] and set the bonus to be ${\mathcal{B}}:=\alpha^{-\frac{1}{2}}\sqrt{K(x_{t},x_{t})-\bm{k}_{x_{t}}^{\top}(\bm{K}+\alpha I)^{-1}\bm{k}_{x_{t}}}$, where $\bm{k}_{x_{t}}:=[K(X,x_{t})]_{(X,R)\in{\mathcal{D}}_{a}}^{\top}\in{\mathbb{R}}^{|{\mathcal{D}}_{a}|}$. We test this PromptWise extension with the Radial Basis Function (RBF) kernel and $\alpha=1$, which we denote by PrompWise-RBF, on both the code completion and code translation tasks. We summarize the results in Figure 27-29. Results show that PromptWise-RBF can perform better on the code completion task (Figure 27). It is also observed that PromptWise-RBF tends to pay lower costs than the original PromptWise algorithm.
Figure 27: Performance of PromptWise with RBF-based kernelized logistic regression (PromptWise-RBF). Results are reported for code completion on BigCodeBench (Task 2). Results are averaged over 20 trials.
Figure 28: Performance of PromptWise with RBF-based kernelized logistic regression (PromptWise-RBF). Results are reported for code completion on the HumanEval benchmark (Task 2). Results are averaged over 20 trials.
Figure 29: Performance of PromptWise with RBF-based kernelized logistic regression (PromptWise-RBF). Results are reported for code translation on the HumanEval-X benchmark (Task 3). Results are averaged over 20 trials.
C.6 Results on Chess Puzzle-Solving
We provide numerical results on solving chess puzzles. In this task, we prompt LLMs to solve chess puzzles in the Lichess puzzle database [^7] https://database.lichess.org/#puzzles. Specifically, the model is provided with the initial board in Forsyth-Edwards Notation (FEN) and is asked to predict the best (legal) move. A prompt example is given in Listing LABEL:lst:chess-puzzle-prompt. In the experiments, we use RoBERTa embedding of the prompt as the input context to the PromptWise algorithm and PAK-UCB baseline. We test PromptWise and the baseline methods on five LLMs: GPT-4o, GPT-4o-mini, Claude-3.7-Sonnet, Qwen-Plus, and Deepseek-Chat. The puzzles are randomly selected from the first $200$ instances in the llm-chess-puzzles Github repository [^8] https://github.com/kagisearch/llm-chess-puzzles. The results are summarized in Figure 30.
Figure 30: Chess Puzzle-Solving (Appendix C.6): GPT-4o, GPT-4o-mini, Claude-3.7-Sonnet, Qwen-Plus, and Deepseek-Chat. Results are averaged over $20$ trials.